# Robust Amortized Bayesian Inference with Self-Consistency Losses on Unlabeled Data

### v4

@@ F.1 Model description
[…]
@@
[…]
F.2 Network architecture and training
For the NPEs $q(\theta|y_{i,t})$, we use a neural spline flow [8] with 10 coupling layers of 256 units each utilizing ReLU activation functions, L2 weight regularization with factor $\gamma=10^{-}3$, 5% dropout and a multivariate unit Gaussian latent space. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. For the summary network, we use a long short-term memory layer with 100 output dimensions followed by a sequence of dense layers with output dimensions of 400, 200, 100, and 50, respectively. The inference and summary network are jointly trained using the Adam optimizer with a batch size of 256 for 100 epochs and a fixed learning rate of $5\times 10^{-4}$, followed by 100 epochs with a fixed learning rate of $5\times 10^{-5}$ and a final run of 100 epochs with a learning rate of $5\times 10^{-6}$.
Appendix G Comprehensive results of the neuron activation case study
@@
Appendix H Detailed setup of the MNIST image denoising case study
We implement the jointly amortized posterior and likelihood networks [37] using two normalizing flows with fully connected affine-coupling layers that operate on the flattened 784-pixel vectors. We use the same network architecture that was used by [37]. As both $\theta$ and $x$ are images whose intrinsic dimensionality is significantly lower than their raw pixel count, we use identical 4-layer convolutional neural networks as summary networks for both posterior and likelihood networks. These summary networks terminate in a global average-pooling layer to produce a 128-dimensional summary of the original or blurred image, respectively. The posterior network itself is implemented as a conditional invertible neural network (cINN) consisting of 12 conditional affine-coupling layers; each coupling layer embeds its conditional information via an internal fully connected network with a single hidden layer of 512 units and ReLU activations. The likelihood network adopts exactly the same conditional-coupling architecture. In both cases, we employ a multivariate Student-T distribution in the latent space [1, 37], which enables more stable maximum-likelihood training at elevated learning rates.
@@ Appendix H Detailed setup of the MNIST image denoising case study
For self-consistency loss, the MNIST test set of digit “0” was divided into two subsets comprising 400 and 580 images respectively. The subset with 400 images was used for training self-consistency loss. A Gaussian blur with $\text{PSF}=1.0$ was applied to these images to generate observations ($x^{*}$). No prior blur was applied to these images. This represents a prior misspecification scenario as the simulated images used to train NPLE were already blurred before applying the noisy camera while the MNIST images used for inference do not have a prior blur. This misspecification scenario depicts the effectiveness of utilising self-consistency loss to overcome prior misspecification. The self-consistency loss was activated at epoch 21, with its weight linearly ramped from zero to one by epoch 40. Training was performed using minibatches of 16 images, and 32 consistency samples were drawn to estimate the variance. The inference was performed on the other held-out subset comprising 580 images and all the results in Figures 5, 14 and 15 use the MNIST images from this subset.
Appendix I Comprehensive results of the MNIST image denoising case study
[…]
[0.5ex]0.4pt3mm Refer to caption Refer to caption
Figure 14: More examples of denoising results for MNIST images of digit “0” in the held-out test set. The first row shows ten randomly selected MNIST images ($\theta$), the second row depicts the same images after applying the Gaussian blur ($x$), third and fourth rows depict the mean and standard deviation of 500 posterior samples estimated from the corresponding blurry observations using NPLE + SC based model, and the fifth and sixth rows depict the mean and standard deviation of 500 posterior samples from model based on NPLE only. Incorporating self-consistency loss significantly improves denoising as the means of reconstructed unblurred image are smoother, less-pixelated and better resemble the ground truth. The darker regions in the standard deviation show the regions of higher variability in the outputs. The standard deviation maps of NPLE + SC based approach are far more coherent showing high variability only along the inner and outer edges.
@@ Appendix I Comprehensive results of the MNIST image denoising case study
Appendix J Computational resources for experiments
@@
    MNIST image denoising: The experiments were run on a high-performance compute cluster using a GPU-equipped compute node featuring a single NVIDIA Tesla P100-PCIE with 12 GB of dedicated HBM2 memory, paired with 16 GB of system RAM. The training for the longest experiment took $\sim 100$ minutes.

### v5

@@ G.1 Model description
where $g_{\text{Na}},g_{\text{K}},g_{\text{M}}$ denote the maximum conductances (in $\text{mS/cm}^{2}$) of sodium and two different types of potassium channels; $E_{\text{Na}},E_{\text{K}},E_{\text{leak}}$ are the sodium, potassium, and leak reversal potentials (in mV) for sodium, potassium, and leak currents; $C_{m}$ is the membrane capacitance (in $\mu\text{F}/\text{cm}^{2}$). The membrane voltage $V_{m}(t)$ at time $t$ is obtained by solving a set of five ordinary differential equations. Here, $X\sim\text{LogNormal}(\mu,\sigma^{2})$ denotes a log-normal distribution with $\log(X)\sim\text{Normal}(\mu,\sigma^{2})$.
@@
Numerical integration of the system is performed using a fixed-step Euler method over a time window $[0,60]$ with time step $\Delta t=0.01$. Voltage traces $V_{m}(t)$ are downsampled to every 30th observation, and 200-dimensional time series are simulated according to $y_{i,t}\sim\text{Student-t}(V_{m}(t),0.1^{2},\text{df=10})$.
G.2 Network architecture and training
For the NPEs $q(\theta|y_{i,t})$, we use a neural spline flow (Durkan et al., 2019) with 10 coupling layers of 256 units each utilizing ReLU activation functions, L2 weight regularization with factor $\gamma=10^{-}3$, 5% dropout and a multivariate unit Gaussian latent space. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. For the summary network, we use a long short-term memory layer with 100 output dimensions followed by a sequence of dense layers with output dimensions of 400, 200, 100, and 50, respectively. The inference and summary network are jointly trained using the Adam optimizer with a batch size of 256 for 100 epochs and a fixed learning rate of $5\times 10^{-4}$, followed by 100 epochs with a fixed learning rate of $5\times 10^{-5}$ and a final run of 100 epochs with a learning rate of $5\times 10^{-6}$.
Appendix H Comprehensive results of the neuron activation case study
@@
Appendix I Detailed setup of the MNIST image denoising case study
We implement the jointly amortized posterior and likelihood networks (Radev et al., 2023) using two normalizing flows with fully connected affine‐coupling layers that operate on the flattened 784‐pixel vectors. We use the same network architecture that was used by Radev et al. (2023). As both $\theta$ and $x$ are images whose intrinsic dimensionality is significantly lower than their raw pixel count, we use identical 4-layer convolutional neural networks as summary networks for both posterior and likelihood networks. These summary networks terminate in a global average‐pooling layer to produce a 128-dimensional summary of the original or blurred image, respectively. The posterior network itself is implemented as a conditional invertible neural network (cINN) consisting of 12 conditional affine‐coupling layers; each coupling layer embeds its conditional information via an internal fully connected network with a single hidden layer of 512 units and ReLU activations. The likelihood network adopts exactly the same conditional‐coupling architecture. In both cases, we employ a multivariate Student-T distribution in the latent space (Alexanderson and Henter, 2020; Radev et al., 2023), which enables more stable maximum‐likelihood training at elevated learning rates.
@@ Appendix I Detailed setup of the MNIST image denoising case study
For self-consistency loss, the MNIST test set of digit “0” was divided into two subsets comprising 400 and 580 images respectively. The subset with 400 images was used for training self-consistency loss. A Gaussian blur with $\text{PSF}=1.0$ was applied to these images to generate observations ($x^{*}$). No prior blur was applied to these images. This represents a prior misspecification scenario as the simulated images used to train NPLE were already blurred before applying the noisy camera while the MNIST images used for inference do not have a prior blur. This misspecification scenario depicts the effectiveness of utilising self-consistency loss to overcome prior misspecification. The self-consistency loss was activated at epoch 21, with its weight linearly ramped from zero to one by epoch 40. Training was performed using minibatches of 16 images, and 32 consistency samples were drawn to estimate the variance. The inference was performed on the other held-out subset comprising 580 images and all the results in Figures 5, 16 and 17 use the MNIST images from this subset.
Appendix J Comprehensive results of the MNIST image denoising case study
Figure 16: Examples of denoising results for MNIST images of digit “0” in the held-out test set. The first row shows ten randomly selected MNIST images ($\theta$), the second row depicts the same images after applying the Gaussian blur ($x$), third and fourth rows depict the mean and standard deviation of 500 posterior samples estimated from the corresponding blurry observations using NPLE + SC based model, and the fifth and sixth rows depict the mean and standard deviation of 500 posterior samples from model based on NPLE only. Incorporating self-consistency loss significantly improves denoising as the means of reconstructed unblurred image are smoother, less-pixelated and better resemble the ground truth. The darker regions in the standard deviation show the regions of higher variability in the outputs. The standard deviation maps of NPLE + SC based approach are far more coherent showing high variability only along the edges.
@@
Appendix K Computational resources for experiments
@@
    MNIST image denoising: The experiments were run on a high-performance compute cluster using a GPU‐equipped compute node featuring a single NVIDIA Tesla P100-PCIE with 12 GB of dedicated HBM2 memory, paired with 16 GB of system RAM. The training for the longest experiment took $\sim 100$ minutes.